Operations Report – December 17th

Crew 201 Operations Report 17-Dec-2018


Name of person filing report: César Augusto Serrano Baza

Non-nominal systems: Generator Autostart

Notes on non-nominal systems: None

Generator (hours run): 5:30 pm last night to around 9 am today. Today started at 4:20 pm.

Solar— SOC 64%

Diesel Reading – 88%

Propane Reading – 20%

Ethanol Free Gasoline – N/A gallons.

Water (Axillary tank) – 0 gallons

Water (static tank) – High Level, around 535 -gallons

Axillary to Static tank transfer– no

Gallons transferred: 0

Water in GreenHab – High level (90%)

Water (loft) – Static to Loft Pump used – yes

Water Meter: 6

Toilet tank emptied: YES

Deimos rover used: no, still not functional

Hours: 0

Beginning charge: –

Ending charge: –

Currently charging: no

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Not used

Hours: 63

Beginning charge:68%

Ending charge: 68%

Currently charging: NOT CHARGING

Opportunity rover used: Not used

Hours: 43.9

Beginning charge: 75%

Ending charge: 75%

Currently charging: NOT CHARGING

Curiosity rover used: Used for EVA Training. 1 couple each 25 minutes

Hours: 59.6

Beginning charge: 100%

Ending charge: 80%

Currently charging: YES
Notes on rovers: Curiosity Rover was used today for the EVA training for the whole crew. The rovers Opportunity and Spirit were not charging during the day, so we could not use them.

# Hours the ATVs were Used today: 0

Notes on ATVs: ATVs were taken out of active service and are now parked away from the station.

HabCar used and why, where? Not used.

CrewCar used and why, where? Not used.

General notes and comments: We did a Training EVA in the place called marble ritual

Summary of internet: Nothing to report.

Summary of suits and radios: We used the Suits (8, 9, 10 and 4) for the EVA training. They were cleaned by the inside after using them and charging.

Summary of Hab operations: Nothing to report.

Summary of GreenHab operations: Cool down with an airflow to keep the temperature lower than 70° Fahrenheit. Director will bring new sticky pads.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Director reset the circuit breakers. The RAM still not charging the rovers

Summary of any observatory issues: Musk Solar observatory – Nominal. Robotic Observatory – Nothing to report.

Summary of health and safety issues: The smoke detectors and CO2 has been activated in every place, except in the musk observatory due to there are not any of them. In the science lab, the CO2 detectors were activated.

Questions, concerns, and requests to Mission Support: We asked the Director for technical support for the Rovers. At the time of the report, Opportunity and Spirit are not charging and the cause still unknown.

Copyright © The Mars Society. All rights reserved. | Main Site